# Swin2-MoSE: A New Single Image Super-Resolution Model for Remote Sensing

### v1

@@ 3.2 Mixture of Experts
Conditional computation paradigms have become popular in deep learning because they allow to scale model capacity, particularly in response to large datasets, maintaining a fixed compute budget. Mixture of Experts (MoE) architecture uses this approach to implement sparsity, improving efficiency and performance by only exploiting necessary parts of the model for a given input. Here, the degree of sparsity defines the ratio between the number of blocks that are actually used to compute the output, compared to the total available. In practical terms, the concepts of “sparse parameter count” (SPC) and “active parameter count” (APC) [48] are decoupled, where the first indicates the model total parameter count, and the second indicates the number of parameters used for processing an individual input. This has been achieved by allowing models to selectively activate only the parts of the network that are relevant to the task at hand, rather than activating the entire network for every input. A comprehensive investigation was done firstly in the domain of NLP, scaling a LSTM model up to 137B parameters [15]. The concept of conditional computation has been even more explored in recent years, culminating in the development of models with a staggering 1.6T parameters [46].
@@
Our model uses as building block the Swin2SR [14] Transformer block, which represents the state-of-the-art model in the super-resolution field. As shown in Figure 2(a), similarly to Switch Transformer [46], SwinV2-MoE [45] and V-MoE [47], we decided to replace the MLP of the Transformer block with our MoE-SM module. We replaced all the MLP layers in the blocks, because we believe that MoE-SM modules can significantly contribute to all processing levels, without adding excessive overhead.
Our MoE-SM (see Figure 2(b)) is designed as a Sparsely-Gated Mixture-of-Experts [15] but, differently from the original version, it is designed with an advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts. Our SM is composed by a 2D convolutional layer with the number of input channels equals to the number of active experts per batch element and output channels of 1 (with a kernel size of 3, padding 1). In this way, we implement an efficient and learnable weighted sum of the input as opposed to a simple and not learnable weighted sum, employed in previous works.
Finally, a trainable Gating Network is used to select the most appropriate experts. Additionally, our MoE-SM works with a new per-example strategy instead of a usually used per-token one. More in detail, the GN redirects each input example in the batch to a subset of experts, rather than distributing individual tokens. Unlike the authors of the paper [15], we do not use the noise term within the gating network, but, instead, we use a more simple Top-K Gating mechanism. We chose this solution because we noted that the learned noise continues to grow during the training, destabilizing the learning process of the Gating Network. The formula for the Top-K Gating Network is presented below.
@@
As in the original implementation [15], we use a loss inside the MoE-SM to promote expert balancing utilization. A metric called is defined to evaluate the relative significance of an expert within a given set of examples.
@@ 3.3 Positional Encoding
The self-attention operation in Vision Transformers is permutation-invariant, leading to the loss of important positional information in 2D images. To address this issue, various positional encoding mechanisms have been incorporated to preserve spatial relationships between image regions. Indeed, the most promising ones are the methods that introduce per-head bias, such as RPE [17] and log-spaced CPB [29], and those which introduce a per-channel bias, such as LePE [16]. To the best of our knowledge, apart for APE and RPE inside the Swin Transformer [17], there has been no prior attempt to analyze how the most popular positional encodings interact with each other when used together.
@@ 3.4 Losses
Two of the most used metrics to evaluate the super-resolved image are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) metrics. However, PSNR does not always provide the correct value between overly smoothed image and a sharp one[49], and its values can vary significantly even for nearly indistinguishable images. Similarly, while SSIM is capable of evaluating brightness, contrast, and structure, it still falls short in reliably quantifying perceptual quality [23]. Whenever these metrics are employed as loss functions during the training process, it is evident that their limitations will be reflected in the super-resolved images generated by the model. In other words, the weaknesses of these metrics will be propagated to the output images, potentially resulting in sub-optimal performance.
@@
SSIM Loss. Structural Similarity Index (SSIM) [50] is a metric used to measure the similarity between two images too. The score is calculated by comparing the luminance, contrast, and structural information of two images. The equation for calculating the SSIM value for the s-th spectral channel () is the following:
@@ 4 Experimental Results
Tasks and Datasets. We performed several experiments on different datasets. In particular, for the ablations, we tested the scale on dataset. To directly assess the performance of SISR models, we used both [18] and OLI2MSI [19]. Both datasets consist of pairs of LR and HR images obtained from different satellites, each with distinct resolutions. Finally, we assess the indirect impact of super-resolution image quality in a practical application, namely semantic segmentation, using the SeasoNet [8] dataset. Below, a more detailed description of the chosen datasets can be found.
@@ 4.1 Training Losses
| Losses | Performance | |||||
| NCC | SSIM | MSE | NCC | SSIM | PSNR | |
| 1 | ✓ | 0.9550 | 0.5804 | 16.4503 | ||
| 2 | ✓ | 0.9565 | 0.9847 | 45.5427 | ||
| 3 | ✓ | 0.9546 | 0.9828 | 45.4759 | ||
| 4 | ✓ | ✓ | 0.9572 | 0.9841 | 45.6986 | |
| 5 | ✓ | ✓ | 0.9549 | 0.9828 | 45.5163 | |
| 6 | ✓ | ✓ | ✓ | 0.9555 | 0.9833 | 45.5542 |
@@ 4.2 Positional Encoding
| Positional Encoding | Performance | ||||
| RPE | log CPB | LePE | SSIM | PSNR | |
| 1 | ✓ | 0.9841 | 45.5855 | ||
| 2 | ✓ | 0.9841 | 45.6986 | ||
| 3 | ✓ | 0.9843 | 45.7278 | ||
| 4 | ✓ | ✓ | 0.9845 | 45.8046 | |
| 5 | ✓ | ✓ | 0.9847 | 45.8539 | |
| 6 | ✓ | ✓ | 0.9843 | 45.6945 | |
| 7 | ✓ | ✓ | ✓ | 0.9846 | 45.8185 |
@@ 4.3 Mixture of Experts
MLP # params Performance Latency Arch SM APC SPC SSIM PSNR (s) 1 MLP 32’670 32’670 0.9845 45.8046 0.179 2 MoE 8/2 32’760 132’480 0.9845 45.8647 0.202 3 MoE 8/2 ✓ 32’779 132’499 0.9849 45.9272 0.212
The aim of this last ablation study is to evaluate the ability of MoE-SM to rival against the MLP (row 3 vs 1 of Table 3). To choose the active and sparse number of experts, we followed the choice made by the authors of Mixtral [48] of 8 experts in total but only 2 of them active for each example in input. To be as fair as possible in the comparison with MLP, each expert has half the number of parameters of an MLP. In that way, the number of active weights involved in each computation is almost the same.
Table 3 shows the comparative results. Baseline model (row 1) is the best model from the previous experiment (row 5 on Table 2). Regarding our MoE without SM (row 2), the SSIM value obtained is equal to the one obtained from the MLP, however PSNR increased by . Instead, our MoE with SM (row 3) surpasses MLP by for SSIM and by of PSNR. On the negative side, the replacement of MLP with MoE slightly increases the average latency, due to the non-optimized sequential execution of the two experts on a single GPU.
@@ 4.4 SISR Experiments
Quantitative comparison. We provide a performance comparison of Swin-MoSE with state-of-the-art SR methods on three upscaling factors. For upscaling, we used dataset: B2, B3, B4, B8 sentinel2 and B3, B4, B7, B11 bands, 10m and 5m resolution, respectively. For upscaling, we used OLI2MSI dataset: blue, green, red from Landsat-8 and B2, B3, B4 from Sentinel2 bands, 30m and 10m resolution, respectively. For upscaling, we used dataset: B5, B6, B7, B8A sentinel2 and B8, B9, B10, B11 bands, 20m and 5m resolution, respectively.
Model OLI2MSI SSIM PSNR SSIM PSNR SSIM PSNR Bicubic 0.9883 45.5588 0.9768 42.1835 0.9674 42.0499 SwinIR [27] 0.9938 48.7064 0.9860 43.7482 0.9825 45.3460 Swinfir [28] 0.9940 48.8532 0.9863 44.4829 0.9830 45.5500 Swin2SR [14] 0.9942 49.0467 0.9881 44.9614 0.9828 45.4759 Swin2SR-MoSE (ours) 0.9948 49.4784 0.9912 45.9194 0.9849 45.9272
The quantitative results shown on Table 4 demonstrate the superiority of our Swin2-MoSE compared to competing SOTA models, improving PSNR by up to dB.
Qualitative comparison. Figures 5, 6,7, and 8 illustrate the qualitative comparison of the considered super-resolution (SR) methods and the proposed approach in this study, for the task of SISR on dataset. Figures 9, 10 and 11 illustrate the qualitative comparison for the task of SISR on OLI2MSI dataset. The red boxes overlaid on the figures highlight the reconstructed regions that are being compared. From the images we can observe the better quality and less distorted images generated from our model compared to SOTA models.
@@ 4.5 Semantic Segmentation Experiments
Finally, we want to demonstrate the effectiveness of the proposed model trained for image super-resolution in improving the performance of a different task, namely semantic segmentation. In order to simplify the procedure, we take into account as input only the 10 m bands of Sentinel-2: B2, B3, B4 and B8. In this way, we can use our Swin2-MoSE, previously trained on SISR task on dataset and frozen, as feature extractor for the semantic segmentation model, Foreground-Aware Relation Network (FarSeg) model [51] in this case. Our goal is to prove that our model is able to provide enriched features the segmentation model, increasing its classification capabilities.
@@ 5 Conclusions
In this work, we presented Swin V2 Mixture of Super-resolution Experts (Swin2-MoSE), a novel End-to-End Single-Image Super-Resolution model tailored on remote sensing that introduced several novelties w.r.t. the state of the art. Firstly, we introduced an an enhanced Sparsely-Gated Mixture-of-Experts (MoE-SM) inside the Transformer architecture blocks increasing the model performance. MoE-SM was designed with a new advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one. Secondly, after an analysis of how positional encodings interact with each other, a per-channel bias (Locally-enhanced Positional Encoding, LePE [16]) and a per-head bias (Relative Position Encoding, RPE [17]) were integrated in the architecture. Thirdly, we proposed a combination of Normalized-Cross-Correlation (NCC) and Structural Similarity Index Measure (SSIM) losses for training, in order to avoid typical limitations of the Mean Square Error (MSE) loss. Finally, several experiments performed on multiple super-resolution satellite images datasets proved that our system is able to surpass the current SOTA both in SSIM and PSNR metrics. Our Swin2-MoSE outperforms competitor models by up to dB (PSNR) and (SSIM) on task of upscaling resolution of factor ( dataset), (OLI2MSI dataset) and ( dataset), respectively. Furthermore, an additional experiment on semantic segmentation showed the capability of our model to enhance the performance of different tasks other than SR alone.

### v2

@@ 3.2 Mixture of Experts
Conditional computation paradigms have become popular in deep learning because they allow to scale model capacity, particularly in response to large datasets, maintaining a fixed compute budget. Mixture of Experts (MoE) architecture uses this approach to implement sparsity, improving efficiency and performance by only exploiting necessary parts of the model for a given input. Here, the degree of sparsity defines the ratio between the number of blocks that are actually used to compute the output, compared to the total available. In practical terms, the concepts of “sparse parameter count” (SPC) and “active parameter count” (APC) 47 are decoupled, where the first indicates the model total parameter count, and the second indicates the number of parameters used for processing an individual input. This has been achieved by allowing models to selectively activate only the parts of the network that are relevant to the task at hand, rather than activating the entire network for every input. A comprehensive investigation was done firstly in the domain of NLP, scaling a LSTM model up to 137B parameters 15. The concept of conditional computation has been even more explored in recent years, culminating in the development of models with a staggering 1.6T parameters 46.
@@
Our model uses as building block the Swin2SR 14 Transformer block, which represents the state-of-the-art model in the super-resolution field. As shown in Figure 2(a), similarly to Switch Transformer 46, SwinV2-MoE 45 and V-MoE 28, we decided to replace the MLP of the Transformer block with our MoE-SM module. We replaced all the MLP layers in the blocks, because we believe that MoE-SM modules can significantly contribute to all processing levels, without adding excessive overhead.
Our MoE-SM (see Figure 2(b)) is designed as a Sparsely-Gated Mixture-of-Experts 15 but, differently from the original version, it is designed with an advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts. Our SM is composed by a 2D convolutional layer with the number of input channels equals to the number of active experts per batch element and output channels of 1 (with a kernel size of 3, padding 1). In this way, we implement an efficient and learnable weighted sum of the input as opposed to a simple and not learnable weighted sum, employed in previous works.
Finally, a trainable Gating Network is used to select the most appropriate experts. Additionally, our MoE-SM works with a new per-example strategy instead of a usually used per-token one. More in detail, the GN redirects each input example in the batch to a subset of experts, rather than distributing individual tokens. Unlike the authors of the paper 15, we do not use the noise term within the gating network, but, instead, we use a more simple Top-K Gating mechanism. We chose this solution because we noted that the learned noise continues to grow during the training, destabilizing the learning process of the Gating Network. The formula for the Top-K Gating Network is presented below.
@@
As in the original implementation 15, we use a loss inside the MoE-SM to promote expert balancing utilization. A metric called is defined to evaluate the relative significance of an expert within a given set of examples.
@@ 3.3 Positional Encoding
The self-attention operation in Vision Transformers is permutation-invariant, leading to the loss of important positional information in 2D images. To address this issue, various positional encoding mechanisms have been incorporated to preserve spatial relationships between image regions. Indeed, the most promising ones are the methods that introduce per-head bias, such as RPE 17 and log-spaced CPB 27, and those which introduce a per-channel bias, such as LePE 16. To the best of our knowledge, apart for APE and RPE inside the Swin Transformer 17, there has been no prior attempt to analyze how the most popular positional encodings interact with each other when used together.
@@ 3.4 Losses
Two of the most used metrics to evaluate the super-resolved image are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) metrics. However, PSNR does not always provide the correct value between overly smoothed image and a sharp one48, and its values can vary significantly even for nearly indistinguishable images. Similarly, while SSIM is capable of evaluating brightness, contrast, and structure, it still falls short in reliably quantifying perceptual quality 21. Whenever these metrics are employed as loss functions during the training process, it is evident that their limitations will be reflected in the super-resolved images generated by the model. In other words, the weaknesses of these metrics will be propagated to the output images, potentially resulting in sub-optimal performance.
@@
SSIM Loss. Structural Similarity Index (SSIM) 49 is a metric used to measure the similarity between two images too. The score is calculated by comparing the luminance, contrast, and structural information of two images. The equation for calculating the SSIM value for the s-th spectral channel () is the following:
@@ 4 Experimental Results
Tasks and Datasets. We performed several experiments on different datasets. In particular, for the ablations, we tested the scale on dataset. To directly assess the performance of SISR models, we used both 50 and OLI2MSI 51. Both datasets consist of pairs of LR and HR images obtained from different satellites, each with distinct resolutions. Finally, we assess the indirect impact of super-resolution image quality in a practical application, namely semantic segmentation, using the SeasoNet 8 dataset. Below, a more detailed description of the chosen datasets can be found.
@@ 4.1 Training Losses
Losses Performance NCC SSIM MSE NCC SSIM PSNR 1 ✓ 0.9550 0.5804 16.4503 2 ✓ 0.9565 0.9847 45.5427 3 ✓ 0.9546 0.9828 45.4759 4 ✓ ✓ 0.9572 0.9841 45.6986 5 ✓ ✓ 0.9549 0.9828 45.5163 6 ✓ ✓ ✓ 0.9555 0.9833 45.5542
@@ 4.2 Positional Encoding
Additionally, to demonstrate the validity of our choice, a more detailed analysis of models with different combinations of positional encoding techniques is proposed in Figure 5. More in detail, the figure presents RPE (, where is the batch size and is the number of patches in a window) and LePE (with shape , where is the number of channels) bias matrices after training. In particular, subfigures (a) and (b) show the matrices of the 6 heads in the first block using the RPE bias without (a) and in combination with (b) LePE bias. Indeed, the behaviour of RPE is not negatively affected by LePE. On the contrary, the patterns appear consolidated. Finally, Figure 5 (c) presents the histograms of the mean value per-channel of the LePE bias in the first block without RPE (left) and with RPE (right) bias. Indeed, the distribution of values using LePE alone or in combination with RPE remain for the most part unchanged indicating that the behavior of the LePE is not negatively influenced by the presence of the RPE.
Positional Encoding Performance RPE log CPB LePE SSIM PSNR 1 ✓ 0.9841 45.5855 2 ✓ 0.9841 45.6986 3 ✓ 0.9843 45.7278 4 ✓ ✓ 0.9845 45.8046 5 ✓ ✓ 0.9847 45.8539 6 ✓ ✓ 0.9843 45.6945 7 ✓ ✓ ✓ 0.9846 45.8185
@@ 4.3 Mixture of Experts
MLP # params Performance Latency Model Model Arch SM APC SPC SSIM PSNR (s) params FLOPs 1 MLP 32’670 32’670 0.9847 45.8539 0.194 2.20M 135.54G 2 MoE 8/2 33’480 131’760 0.9845 45.8647 0.223 4.59M 135.68G 3 MoE 8/2 ✓ 33’499 131’779 0.9849 45.9272 0.229 4.59M 138.30G 4 MoE 8/1 ✓ 34’120 262’810 0.9837 45.4847 0.230 7.72M 136.96G 5 MoE 6/2 ✓ 33’319 98’839 0.9844 45.8328 0.230 3.80M 138.30G 6 MoE 8/2 ✓ 33’499 131’779 0.9849 45.9272 0.229 4.59M 138.30G 7 MoE 10/2 ✓ 33’679 164’719 0.9846 45.8913 0.230 5.39M 138.30G
The aim of this last ablation study is to evaluate the ability of MoE-SM to rival against the MLP (row 3 vs 1 of Table 3). To choose the active and sparse number of experts, we followed the choice made by the authors of Mixtral 47 of 8 experts in total but only 2 of them active for each example in input. To be as fair as possible in the comparison with MLP, each expert has half the number of parameters of an MLP. In that way, the number of active weights involved in each computation is almost the same. The 8/2 choice is subsequently validated by a further ablation where the total number of experts available and the number of experts actually used during the forward pass were varied.
Table 3 shows the comparative results. Baseline model (row 1) is the best model from the previous experiment (row 5 on Table 2), which uses MLP. Regarding our MoE without SM (row 2), the SSIM and PSNR have some fluctuation compared to MLP version (row 1). Instead, our MoE with SM (row 3) surpasses MLP by for SSIM and by of PSNR. On the negative side, the replacement of MLP with MoE slightly increases the average latency, due to the non-optimized sequential execution of the two experts on a single GPU.
With regard to the second ablation study (from row 4 to row 7), reducing the number of active experts (row 4 vs. row 6) decreases the performance, highlighting the important role played by our SM module. Also decreasing (row 5 vs. row 6) or increasing (row 7 vs. row 6) the number of total experts available does not help. In our hypothesis, the choice of the number of total experts could also be linked to the amount of data available during training. For this reason, the strategy of further increasing the number of total experts could be fruitful if the dataset were larger.
@@ 4.4 SISR Experiments
Quantitative comparison. We provide a performance comparison of Swin-MoSE with state-of-the-art SR methods on three upscaling factors. For upscaling, we used dataset: B2, B3, B4, B8 sentinel2 and B3, B4, B7, B11 bands, 10 m and 5 m resolution, respectively. For upscaling, we used OLI2MSI dataset: blue, green, red from Landsat-8 and B2, B3, B4 from Sentinel2 bands, 30 m and 10 m resolution, respectively. For upscaling, we used dataset: B5, B6, B7, B8A sentinel2 and B8, B9, B10, B11 bands, 20 m and 5 m resolution, respectively. For others method, we supposed to use the usual MSE loss for training, compared to us, who use SSIM and NCC together as described in our theory section.
Model Params SSIM PSNR GiB IT/s APC SPC FLOPs Bicubic 0.9883 45.5588 SwinIR 25 0.9938 48.7064 2.37 1.32 2.10M 2.10M 515.90G Swinfir 26 0.9940 48.8532 3.12 1.28 2.52M 2.52M 622.42G Swin2SR 14 0.9942 49.0467 3.12 1.18 2.11M 2.11M 526.78G DAT 30 0.9946 49.4699 1.46 1.95 2.14M 2.14M 547.22G Swin2-MoSE (ours) 0.9948 49.4883 2.41 1.16 2.20M 4.56M 543.00G OLI2MSI Params SSIM PSNR GiB IT/s APC SPC FLOPs Bicubic 0.9768 42.1835 SwinIR 25 0.9863 44.4806 3.55 0.77 2.11M 2.11M 809.42G Swinfir 26 0.9863 44.4829 4.73 0.76 2.53M 2.53M 975.84G Swin2SR 14 0.9881 44.9614 4.80 0.70 2.12M 2.12M 826.30G DAT 30 0.9886 45.0404 2.10 1.10 2.15M 2.15M 858.35G Swin2-MoSE (ours) 0.9912 45.9414 4.88 0.52 2.20M 4.56M 851.76G Params SSIM PSNR GiB IT/s APC SPC FLOPs Bicubic 0.9674 42.0499 SwinIR 25 0.9825 45.3460 0.84 5.37 2.14M 2.14M 131.53G Swinfir 26 0.9830 45.5500 0.84 5.20 2.56M 2.56M 158.17G Swin2SR 14 0.9828 45.4759 1.03 4.74 2.15M 2.15M 134.40G DAT 30 0.9837 45.8053 0.69 6.73 2.18M 2.18M 139.36G Swin2-MoSE (ours) 0.9849 45.9272 0.86 3.91 2.23M 4.59M 138.30G
The quantitative results shown on Table 4 demonstrate the superiority of our Swin2-MoSE compared to competing SOTA models. It improves PSNR by up to dB compared to compared with models descended from Swin Transformer like ours. Compared to SOTA, it slightly outperforms for the 2x task, but maintains a good gap compared to 3x and 4x tasks. This demonstrates the goodness of our model which behaves significantly better in more complex situations. Additionally, the table presents an evaluation of the computational complexity of our model. Even if our model has an higher complexity in terms of number of sparse parameters due to the MoE, the number of active parameters and floating-point operations (FLOPs) is still comparable to the other state-of-the-art models.
Qualitative comparison. Figures 6, 7, 8, and 9 illustrate the qualitative comparison of the considered super-resolution (SR) methods and the proposed approach in this study, for the task of SISR on dataset. Figures 10, 11 and 12 illustrate the qualitative comparison for the task of SISR on OLI2MSI dataset. The red boxes overlaid on the figures highlight the reconstructed regions that are being compared. From the images we can observe the better quality and less distorted images generated from our model compared to SOTA models.
@@ 4.5 Semantic Segmentation Experiments
Finally, we want to demonstrate the effectiveness of the proposed model trained for image super-resolution in improving the performance of a different task, namely semantic segmentation. In order to simplify the procedure, we take into account as input only the 10 m bands of Sentinel-2: B2, B3, B4 and B8. In this way, we can use our Swin2-MoSE, previously trained on SISR task on dataset and frozen, as feature extractor for the semantic segmentation model, Foreground-Aware Relation Network (FarSeg) model 52 in this case. Our goal is to prove that our model is able to provide enriched features the segmentation model, increasing its classification capabilities.
@@ 5 Conclusions
In this work, we presented Swin V2 Mixture of Super-resolution Experts (Swin2-MoSE), a novel End-to-End Single-Image Super-Resolution model tailored on remote sensing that introduced several novelties w.r.t. the state of the art. Firstly, we introduced an an enhanced Sparsely-Gated Mixture-of-Experts (MoE-SM) inside the Transformer architecture blocks increasing the model performance. MoE-SM was designed with a new advanced merging layer (Smart Merger, SM) in order to merge the output of individual experts, and with a new way to split the work between experts, defining a new per-example strategy instead of the commonly used per-token one. Secondly, after an analysis of how positional encodings interact with each other, a per-channel bias (Locally-enhanced Positional Encoding, LePE 16) and a per-head bias (Relative Position Encoding, RPE 17) were integrated in the architecture. Thirdly, we proposed a combination of Normalized-Cross-Correlation (NCC) and Structural Similarity Index Measure (SSIM) losses for training, in order to avoid typical limitations of the Mean Square Error (MSE) loss. Finally, several experiments performed on multiple super-resolution satellite images datasets proved that our system is able to surpass the current SOTA both in SSIM and PSNR metrics. Our Swin2-MoSE outperforms competitor Swin-based models by up to dB (PSNR) and (SSIM) on task of upscaling resolution of factor ( dataset), (OLI2MSI dataset) and ( dataset), respectively. Compared to SOTA, it achieves better overall performance, especially for more difficult tasks such as 3x and 4x, where it shows great potential. Furthermore, an additional experiment on semantic segmentation showed the capability of our model to enhance the performance of different tasks other than SR alone.